OpenStudy (sanic):

The graph shows the amount of money paid when purchasing bags of caramel corn at the zoo: A graph is shown. The x-axis is labeled from 0 to 9. The y-axis is labeled from 0 to 10. The four points are shown on the graph are on ordered pairs 0, 0 and 1, 3 and 2, 6 and 3, 9. These points are joined by a line. The label on the x-axis is Bags of caramel corn. The title on the y-axis is Total cost in dollars. Write an equation to represent the relationship between the total cost (y) and the number of bags of caramel corn (x). y = one-half x y = 2x y = 3x y = one-third x

OpenStudy (sanic):

Plz help

OpenStudy (sanic):

I'll give medal

OpenStudy (anonymous):

Okay lets work this out. @Sanic

OpenStudy (anonymous):

Its C.

OpenStudy (sanic):

Ok but I'll work it out

OpenStudy (anonymous):

One point is 2,6. y/x or 6/2 is 3. C.) is Y=3x. 3*2 is 6. so the answer is C.)

OpenStudy (anonymous):

1×3=3 2×3=6 3×3=9 i.e. you multiply each number by three

OpenStudy (sanic):

Ok

OpenStudy (anonymous):

Yeah @BrixtonLewis got it

OpenStudy (anonymous):

Have a great day!

OpenStudy (sanic):

Sorry no more help needed